OpenStudy (anonymous):

x^2 + 6x + 12= (x+3)^2 True False

OpenStudy (anonymous):

My guess is False

OpenStudy (anonymous):

@HelpBlahBlahBlah

OpenStudy (anonymous):

simplify it !

OpenStudy (anonymous):

(x+3)^2 = ??

OpenStudy (anonymous):

x+3 times x+3

OpenStudy (anonymous):

which would be x^6?

OpenStudy (anonymous):

no , it has a rule , (a + b )² = (a² + 2ab + b²) apply the rule !

OpenStudy (anonymous):

(x^2 +2x3+3^2?

OpenStudy (anonymous):

(x² + 2(3*x) + 3² ) = ?

OpenStudy (anonymous):

no idea :/

OpenStudy (anonymous):

2(3x) = ? 3² = ?

OpenStudy (anonymous):

Not to be rude. But your not much help, your making it even more confusing :(

OpenStudy (anonymous):

dude , 2times 3 = 6 , 6x 3² = ?